OpenStudy (shy_boy):

Find the Range of f(w)= 2w^2+w-1 and g(w)=w^2-(underoot)w-1

OpenStudy (shy_boy):

\[f(w)=2w ^{2}+w-1 and g(w)=w ^{2}-\sqrt{w-1}\]

OpenStudy (shy_boy):

@Hannahsayshi

OpenStudy (shy_boy):

How to find the Range

OpenStudy (hannahsayshi):

Okay I'm going to tag my friend okay

OpenStudy (shy_boy):

OK :)

OpenStudy (hannahsayshi):

@A_Burning_Masquerade

OpenStudy (a_burning_masquerade):

Give me two minutes tops

OpenStudy (shy_boy):

OK Sir!

OpenStudy (a_burning_masquerade):

Okay when I do the math the answer comes out differently then when I check it. I really don't know because I'm usually not wrong and I don't want to give you the wrong answer. So could @TheSmartOne help out.

OpenStudy (faiqraees):

find x for f'(x)=0. Plug it in f(x) and evaluate the range

OpenStudy (faiqraees):

Can you do it?

OpenStudy (shy_boy):

OK Faiq how am I supposed to do that can u please explain

OpenStudy (faiqraees):

So there are two methods for you question. One is simply drawing a graph to evaluate it out and second is calculus (the cool one). Which method do you opt for?

OpenStudy (shy_boy):

Its OK np @A_Burning_Masquerade Thanks for trying to help me out

OpenStudy (shy_boy):

calculus :)

OpenStudy (faiqraees):

So first you've to understand what the notation implies. f(x) means something described in the terms of x or you can say a function of x f'(x) or \(\large\rm \frac{d}{dx}f(x) \) implies the change in rate of f(x) with respect to x. Are you clear on that?

OpenStudy (shy_boy):

yeah its clear Thanks

OpenStudy (faiqraees):

So \(\large\rm \frac{d}{dx}f(x)\) implies the change in f(x). Now the change can be positive. Meaning for increasing values of x, the output is increasing. Or the change could be negative. Meaning for increasing values of x, the output is decreasing. But a special case comes at 0. When\(\large\rm \frac{d}{dx}f(x)=0\) it means that on that very x coordinate, the \(\large\rm f(x)\) has gone from decreasing to increasing or from increasing to decreasing. Clear on that?

OpenStudy (shy_boy):

yes sir

OpenStudy (faiqraees):

Now notice the graph of \(\large\rm f(x)\)when \(\large\rm \frac{d}{dx}f(x)=0\)  Can you point out by drawing a line which is the lowest point for f(x)?

OpenStudy (shy_boy):